OpenStudy (anonymous):

jen has only $1 bills and $5 dollar bills. she has 7 bills worth the totalof $19. lets X represent the # of $1 bills and y represent the # of $5 bills. which system of linear equations can b used to determine the number of $1 bills????

OpenStudy (anonymous):

I need help on how to get the answer?

OpenStudy (anonymous):

let the number of 1$ bills be x and so be y for 5$ bills we hav x+y=7 and x+5y=19 solve these two and u'll get x=4 and y=3.

OpenStudy (anonymous):

That's not and answer choice. Want me to post the answer choices? I have the right answer my teacher gave it to us. I just need to see it worked out i dont get how she got the answer.

OpenStudy (anonymous):

post the choices...

OpenStudy (anonymous):

hes right

OpenStudy (anonymous):

ok a) {x+y=7} x+5y=19 B) x+y=19 x=5y=29 C) x+y=6 x+5y=19 D) x+y=7 5x=y=19 Those are the choices...I hve the right answer i just need to see how the teacher got it.

OpenStudy (anonymous):

i guess a is correct. and the steps i've shown u there already i u hav doubt in those, ask me specifying wer u have one :)

OpenStudy (anonymous):

yeah the answer is A bur i dont how u or the teacher got it.

OpenStudy (anonymous):

ok wait, in the question we need to find, the number of 1$ bills. the information given to us is that , if u add the total value of the money u hav it is 19$ . and the total number of bills (paper sheets-money note) u have is 7. therefore if i assume number of 1$ bills be x and dat of 5$ be y, then x+y=7.(since the total number of notes u hav is = to individual notes' count) and x+5*y =19(since total money worth is the sum of individual contribution...like if u hav 3 1$ notes and 2 5$ notes...total money u hav is 1+1+1+5+5=3*1+2*5 , which is what i've done) so now solve these two linear equations to get x. this is the best possible explanation i can give :|

OpenStudy (anonymous):

Okie dokie. Thankyou:)